OpenStudy (anonymous):

Write the equation 2x+7y-4=0 in normal form. than find the length of the normal and the angle it makes with the positive x-axis

OpenStudy (anonymous):

what do you mean by normal form?

OpenStudy (anonymous):

do you know when equations are written like this" 2x/squrt69 + 7y/squrt69-4/squrt69=0

OpenStudy (wolf1728):

That recent question has nothing to do with the first right?

OpenStudy (anonymous):

no i mean their on the same work sheet but no

OpenStudy (anonymous):

they're

OpenStudy (wolf1728):

So let's try to solve one at a time huh?

OpenStudy (anonymous):

ok :)

OpenStudy (wolf1728):

7y = 4 - 2x y = 4/7 -2x/7 So slope = -2/7 There's another way to check that

OpenStudy (anonymous):

wait why is slope important?

OpenStudy (anonymous):

hello?

OpenStudy (wolf1728):

Geez, slope is important because it shows the rate of change of a function.

OpenStudy (anonymous):

sorry im worried i missed a bunch of school and my friends notes are not helpful

OpenStudy (anonymous):

so we have to graph it?

OpenStudy (wolf1728):

FORGOT THE X ARRRGGGGHHHHHHHHHHHHH!!!!!!!!!!!

OpenStudy (wolf1728):

y = 4/7 -2x/7 If x = 1 then y = (4/7) -2*(1/7) y = 2/7 y = 0.2857142857 Produces an equation of Y = -0.2857142857 X + 0.5714285714 If we multiply both sides by 7 we get 7y = -2X + 4

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

We have to find the angle now. If you missed a lot of class(es) I'd say you should review that material first. (I could still calculate the angle for you) :-)

OpenStudy (anonymous):

i found the online lesson and now im even more confused ._.

OpenStudy (anonymous):

i can fail one unit ill live... probably

OpenStudy (wolf1728):

Since slope is -2/7 we know it slopes from "high" on the left to "low" on the right. I think the angle would be the arc sine of -2/7 which would be -16.602 Degrees Well how soon do you have to finish this?

OpenStudy (anonymous):

no i figured it out, because i only need normal form ist jsut taking AB and C and putting them into some weird equation, thank you for trying to help :) i just have to have it doen tonight! much thanks

OpenStudy (wolf1728):

I can graph it for you - really not much trouble - computre will do it.

OpenStudy (anonymous):

the angle isnt the biggest deal but i got the angle is =74 and the length of the normal is =4/sqrt53 =.55 units

OpenStudy (anonymous):

but with the angle being 16 the normal length is either 4/squrt13 or 4/squrt9

OpenStudy (anonymous):

could you by chance explain the one that no body answered? it's which is the standard form of the equation p=18 Phi=5pi/4

OpenStudy (wolf1728):

Here's the graph

OpenStudy (anonymous):

thank you

OpenStudy (wolf1728):

I redid it now it's only one line showing (don't know why - I'm getting tired) Here's the data to go with it: X Y -10 24 -9 22 -8 20 -7 18 -6 16 -5 14 -4 12 -3 10 -2 8 -1 6 0 4 1 2 2 0 3 -2 4 -4 5 -6 6 -8 7 -10 8 -12 9 -14 10 -16

OpenStudy (anonymous):

its ok you can sleep i think ive got the hang of it :)

OpenStudy (wolf1728):

I believe I did NOT get the equation for the graph correctly. ARRGGHHHH!!!!!!!!!!!!!!!!!!!!!!!!!!!!! My Fault I'll re-redo it - very close to finishing it

OpenStudy (wolf1728):

Here we go yet again - it has the right slope!!!!!!!!!!!!!!!!!!!!! Data X Y -10.00 3.43 -9.00 3.14 -8.00 2.86 -7.00 2.57 -6.00 2.29 -5.00 2.00 -4.00 1.71 -3.00 1.43 -2.00 1.14 -1.00 0.86 0.00 0.57 1.00 0.29 2.00 0.00 3.00 -0.29 4.00 -0.57 5.00 -0.86 6.00 -1.14 7.00 -1.43 8.00 -1.71 9.00 -2.00 10.00 -2.29

OpenStudy (anonymous):

its ok its ok i got that asnwer right theres no worries hun

OpenStudy (wolf1728):

In the words of Inspector Clousseau "Once again I have failed where others have succeeded." LOL

OpenStudy (anonymous):

i just got stumped on the angle ones because the directions were confuseing but ill get it its ok ;)

OpenStudy (anonymous):

3x-2y-1=0 is 326 degrees 5x=y-12=0 is 11 degrees 4x=3y-4=0 is 37 degrees

OpenStudy (anonymous):

those are the ones i got wrong and their correct answers

OpenStudy (wolf1728):

To get the angle - you'd need the slope which is -2/7 Then look up the arc sine of -2/7 (or -0.2857142857) okay - guess I'll get going

OpenStudy (anonymous):

sweet dreams! btw i got a 100 on the practice quiz just now!

OpenStudy (wolf1728):

okay - well done !!!! must get going see ya